OpenStudy (anonymous):

Given that the pair of triangles is similar, find the length of the side labeled x. 1st triangle is 13.8 & x 2nd triangle is 6.9 & 8.9

OpenStudy (anonymous):

its 45.5

OpenStudy (anonymous):

sorry 17.8

OpenStudy (anonymous):

^ it's 13.8 not 3.8

Directrix (directrix):

Lengths of corresponding sides of similar triangle are in proportion. It is not clear from the problem which side corresponds to which. I'm guessing: (13.8)/(6.9) = (x)/(8.9) where x = 17.8 approx

Directrix (directrix):

@ Risky -- Thanks